Evaluate the integral x / x+1 dx 
1
Expert's answer
2021-09-10T14:37:04-0400
"\int \dfrac{x}{x+1}dx=\int \dfrac{x+1-1}{x+1}dx=\int(1-\dfrac{1}{x+1})dx"

"=x-\ln|x+1|+C"